OpenStudy (jamanorthi):

A bag of tiles contains 5 A's, 3 B's, 2 C's. what is the probability of picking ABC in order in three draws from the bag, without replacement?

OpenStudy (jamanorthi):

someone help plzzzz

OpenStudy (mathstudent55):

For the first pick, what is the probbaility of picking an A?

OpenStudy (jamanorthi):

i don't know how to do that

OpenStudy (mathstudent55):

Ok. The probability of an event happening is the number of ways the event can happen divided by the total number of ways all events can happen.

OpenStudy (mathstudent55):

In this case, all events are picking one of 10 tiles. The total number of all events happening is, therefore, 10. Ok?

OpenStudy (jamanorthi):

is picking A is 5/30

OpenStudy (mathstudent55):

Correct.

OpenStudy (mathstudent55):

Once an A has been picked, how many tiles are left?

OpenStudy (jamanorthi):

okay so i do the same thing for all of them and then just multiply them together the fraction

OpenStudy (mathstudent55):

Yes, but the total number of total events changes each time. Once the A was picked on the first pick, since there is no replacement, there are how many tiles left for the second pick?

OpenStudy (jamanorthi):

29

OpenStudy (jamanorthi):

oh no there will be 25

OpenStudy (mathstudent55):

I think you mean 9. There were 5 + 3 + 2 tiles = 10 tiles. An A was taken out. There are now 10 - 1 = 9 tiles for the second pick.

OpenStudy (jamanorthi):

yeah you are right

OpenStudy (jamanorthi):

then picking A is 5/10 right??

OpenStudy (mathstudent55):

Yes, for the first pick, the probability of picking an A is 5/10. For the second pick, you want a B out of 9 tiles. How many B's are there?

OpenStudy (jamanorthi):

3 B's so 3/9

OpenStudy (mathstudent55):

Correct. Now for the third pick, there are only 8 tiles left. How many C's are there?

OpenStudy (jamanorthi):

2 C's so 2/8

OpenStudy (mathstudent55):

Correct. All these fractions can be simplified. To find the answer, multiply them together.

OpenStudy (jamanorthi):

so is the answer 1/24

OpenStudy (mathstudent55):

5\10 * 3/9 * 2/8 = = 1/2 * 1/3 * 1/4 = 1/24 You are correct.

OpenStudy (jamanorthi):

@ mathstudent55 thank you so much

OpenStudy (mathstudent55):

You are very welcome.